### v2

# LEXA: Legal Case Retrieval via Graph Contrastive Learning with Contextualised LLM Embeddings

Abstract
Legal case retrieval (LCR) is a specialised information retrieval task aimed at identifying relevant cases given a query case. LCR holds pivotal significance in facilitating legal practitioners to locate legal precedents. Existing LCR methods predominantly rely on traditional lexical models or language models; however, they typically overlook the domain-specific structural information embedded in legal documents. Our previous work CaseGNN [casegnn] successfully harnesses text-attributed graphs and graph neural networks to incorporate structural legal information. Nonetheless, three key challenges remain in enhancing the representational capacity of CaseGNN: (1) The under-utilisation of rich edge information in text-attributed case graph (TACG). (2) The insufficiency of training signals for graph contrastive learning. (3) The lack of contextualised legal information in node and edge features. In this paper, the LEXA model, an extension of CaseGNN, is proposed to overcome these limitations by jointly leveraging rich edge information, enhanced training signals, and contextualised embeddings derived from large language models (LLMs). Specifically, an edge-updated graph attention layer (EUGAT) is proposed to comprehensively update node and edge features during graph modelling, resulting in a full utilisation of structural information of legal cases. Moreover, LEXA incorporates a novel graph contrastive learning objective with graph augmentation to provide additional training signals, thereby strengthening the model’s legal comprehension capabilities. What’s more, given the remarkable contextualised understanding capabilities of LLMs for text encoding, LLMs are employed to generate node and edge features for the text-attributed case graph (TACG). Extensive experiments on two benchmark datasets from COLIEE 2022 and COLIEE 2023 demonstrate that LEXA not only significantly improves CaseGNN but also achieves supreme performance compared to state-of-the-art LCR methods. Code has been released on here.
keywords:
Information Retrieval, Legal Case Retrieval, Graph Neural Networks
1 Introduction
Given a query case, legal case retrieval (LCR) aims to find relevant cases in a large case database, and has become an indispensable tool for legal practitioners such as judges and lawyers. Recently, the development of LCR tools has significantly streamlined this process, enabling rapid and accurate retrieval of relevant cases. Moreover, LCR tools also serve as a convenient resource for individuals seeking legal consultation but lacking the financial means to afford expensive legal services. In recent years, LCR approaches have been mostly based on two methodologies: lexical retrieval models and language models (LMs). Traditional lexical retrieval models, including BM25 [BM25], TF-IDF [TF-IDF] and LMIR [LMIR], calculate the similarity score between cases based on term frequency. Language models [Law2Vec, Lawformer, MTFT-BERT, MVCL, BERT-PLI, LEGAL-BERT, SAILER, JOTR, DoSSIER, RPRS, NOWJ, ConversationalAgent, UA@COLIEE2022, CL4LJP] rely on pre-trained models to encode cases into case representations for calculating case similarity. Specifically, due to the intricate structure of legal texts, language model-based LCR methods employ various case similarity comparison strategies, including sentence-level [IOT-Match], paragraph-level [BERT-PLI], and document-level [SAILER] approaches.
However, important legal structural information is neglected in LM-based LCR models, as they rely solely on raw text to generate high-dimensional representations of cases for similarity calculation. The structural information of a legal case captures the relationships among its core elements, such as parties, criminal activities, and evidence, which can help to further improve the understanding capability of the model. Therefore, in our previous work, CaseGNN [casegnn] is proposed to leverage legal structural relationships to improve LCR accuracy. Firstly, the structural information of legal cases is extracted in the form of relational triplets, which are then encoded by LMs to generate node features for the constructed Text-Attributed Case Graph (TACG). With the constructed TACG, a CaseGNN framework is proposed to generate the final case representation. Finally, to train the CaseGNN model, a contrastive learning loss is designed to capture the training signal from both easy and hard negative samples.
Although CaseGNN exploits the legal structural information that is neglected in the lexical models and LMs, three key challenges remain unaddressed, hindering the accuracy and effectiveness of LCR. (1) Underutilisation of rich edge information: CaseGNN only focuses on updating node features during graph modelling, thereby neglecting the rich and informative edge features inherent in the TACG. The edges in the TACG represent relationships between nodes, capturing legal connections between entities, which are crucial for accurately generating case representations via graph neural network models. However, in the GNN layer of CaseGNN, the edge features remain fixed throughout the training process, preventing the model from leveraging updated relational information derived from evolving node features. (2) Insufficiency of training signals: The training process of the CaseGNN model is hindered by a scarcity of training signals, primarily due to the limited availability of labelled legal data. This challenge is further exacerbated by the restricted access to legal sources and the high cost associated with annotating legal datasets. Since the labelling process typically requires legal expertise from professionals such as judges, lawyers, or legal scholars, it incurs substantial time and financial costs. Consequently, the inadequacy of labelled data leads to insufficient training signals for effectively training LCR models. (3) The lack of contextualised legal information in node and edge features: In CaseGNN, the node and edge features are encoded by LMs, which lacks contextualised legal information. This limitation restricts the model’s ability to capture the intricate legal semantics and interrelations embedded in legal texts, thereby affecting the expressiveness of the final case representations generated from case graphs constructed with suboptimal node and edge features.
To address the identified limitations of CaseGNN, an enhanced framework, LEXA, is proposed to extend CaseGNN [casegnn] in a more accurate and effective manner. First, LEXA collaboratively exploits and dynamically updates both node and edge features by proposing a novel Edge-updated Graph Attention Layer to generate more comprehensive and informative case representations. Second, to mitigate the issue of insufficient training signals, graph contrastive learning with graph augmentation is employed, providing additional supervision to enhance the model’s representational capacity. Finally, a top-performing text encoding-oriented large language model (LLM) is leveraged to encode node and edge features, capturing contextualised legal semantics embedded within legal cases. Empirical experiments are conducted on two benchmark datasets, COLIEE 2022 [COLIEE2022] and COLIEE 2023 [COLIEE2023], demonstrating that LEXA not only outperforms the original CaseGNN model but also achieves state-of-the-art performance on the legal case retrieval (LCR) task. Compared with CaseGNN [casegnn], the main contributions of this paper are summarised as follows:
  • A LEXA framework is proposed to address the limitations of CaseGNN and to further improve retrieval accuracy and effectiveness by fully exploiting edge information, providing additional training signals, and leveraging LLMs to incorporate contextualised legal semantics.
  • A GNN layer called Edge-updated Graph Attention Layer (EUGAT) is designed to simultaneously update both node and edge features. This mechanism enables the generation of comprehensive and informative case representations for model training.
  • A graph contrastive learning (GCL) objective is developed, along with a graph augmentation strategy to provide supplementary training signals that enhance the model’s legal understanding capabilities.
  • A top-performing text-embedding LLM, LEXA-8B, is fine-tuned and employed to encode node and edge features, enabling richer contextualised and latent legal semantics to be captured within legal cases.
  • Extensive experiments conducted on the COLIEE 2022 and 2023 benchmark datasets validate the superior performance of LEXA, demonstrating significant improvements over CaseGNN and outperforming existing state-of-the-art LCR approaches.
2 Related Work
In this section, the related work will be reviewed in detail from three aspects that related to this paper, which are legal case retrieval, graph contrastive learning, and text embedding LLM.
2.1 Legal Case Retrieval
Legal Case Retrieval, as a specialised subtask of information retrieval, has seen the development of two main categories of methods: traditional statistical retrieval models and neural language models. Statistical models such as TF-IDF [TF-IDF], BM25 [BM25], and LMIR [LMIR] rely on term frequency and inverse document frequency to compute relevance scores between a query and candidate cases. These approaches are efficient and interpretable but often fall short in capturing semantic similarities in complex legal texts. More recently, language models [DeepCT, BERT, RoBERTa, monot5] have gained prominence in LCR due to their strong language understanding capabilities. Numerous LCR-specific adaptations of LMs have been proposed [Law2Vec, Lawformer, MTFT-BERT, MVCL, LEGAL-BERT, JOTR, DoSSIER, RPRS, NOWJ, UA@COLIEE2022, IOT-Match, Law-Match, LEDsummary, BM25injtct, LeiBi, LEVEN, JNLP@COLIEE2019, CL4LJP, QAjudge, query_conversational_agent, ConversationalAgent], demonstrating substantial improvements over traditional retrieval techniques. Specifically, Law2Vec [Law2Vec] utilises a large number of English legal corpus to train a Word2Vec [word2vec] model to get legal word embeddings. Besides, Lawformer [Lawformer] leverages Chinese legal documents to pre-train a longformer [longformer] based model into a Chinese specific legal language model. MVCL [MVCL] exploits a contrastive learning model with case-view and element-view strategies to enhance the understanding ability of the model and improve the performance of relevant case retrieval task. And CL4LJP [CL4LJP] combines three different levels contrastive learning objectives that includes article-level, charge-level and label-aware level to jointly train the legal judgment prediction model. To tackle the long text problem in legal domain, BERT-PLI [BERT-PLI] divides cases into paragraphs and calculates the similarity between two paragraphs while SAILER [SAILER] directly truncates the case text to cope with the input limit of LM. PromptCase [promptcase] proposes a prompt-based encoding scheme with two important legal facts and legal issues to address the lengthy problem. CaseLink [caselink] utilises the intrinsic case connectivity relationships among cases with graph neural networks to improve the retrieval accuracy.
2.2 Graph Contrastive Learning
Contrastive learning, a self-supervised learning method, focuses on learning the feature representation by bringing the positive samples closer while pushing the negative samples away [nce, NCE4NPLMs]. Contrastive learning is widely used in computer vision task [moco, SimCLR], natural language processing task [SimCSE, CLEAR, COCO-LM], and information retrieval task [PSSL, ANCE] for providing additional training signals to train the model instead of extra annotation.
Recently, contrastive learning is also applied in graph domain to fully leverage the graph structural information and explore the unlabelled graph data. The main difference between contrastive learning in the graph domain and other domains is the augmentation methods of graphs focusing on the aspects of graph structure such as node, edge or subgraph, where other domains have no graph structure. Specifically, there are various graph augmentation methods used for graph contrastive learning. Edge dropping is firstly proposed by Rong et al [DropEdge]to remove a certain portion of edges from the original graph. Similarly, node dropping [GRAND, GraphCL] refers to removing nodes with a certain probability. Besides of the graph structure augmentation such as removing nodes and edges, graph feature masking [BGRL, GraphCL] is also a popular graph augmentation method to perform efficient graph augmenting or corrupting by randomly mask the features of nodes or edges. Considering graph can be divided into subgraph, another graph augmentation approach called subgraph cropping is to crop out subgraph from the original graph then use the subgraph itself or the remaining graph as samples for contrastive learning (e.g.,  [GraphCL, GraphCrop]).
2.3 LLM Embeddings
Large language models (LLMs) have demonstrated outstanding performance in encoding text into high-quality embeddings for a wide range of natural language processing tasks, such as information retrieval and question answering. Their impressive encoding capability arises from the ability to model complex linguistic patterns and contextual dependencies through deep transformer architectures pretrained on large-scale corpora. Various LLM-based embedding models have been proposed to enhance semantic representation across diverse tasks. The recently proposed Multilingual Massive Text Embedding Benchmark (MMTEB) [mmteb] offers a comprehensive evaluation framework for assessing text embedding models across multiple tasks, such as classification, retrieval, and clustering. Legal retrieval is a domain-specific task included in MMTEB, where several state-of-the-art text embedding LLMs have achieved notable performance. voyage-law-2111https://blog.voyageai.com/2024/04/15/domain-specific-embeddings-and-retrieval-legal-edition-voyage-law-2/ ranks first in the MMTEB legal retrieval task by leveraging high-quality legal tokens, and a novel contrastive learning algorithm. E5-Mistral-7B-Instruct[e5mistral] is another top-performing model that utilizes proprietary LLMs to generate diverse synthetic data for training. OpenAI’s text-embedding-3-large [^2] https://platform.openai.com/docs/guides/embeddings is a high-quality general purpose embedding model, though it is primarily optimized for English and not specifically designed for multilingual retrieval tasks such as those in MMTEB. Additionally, Qwen3-Embedding-8B[qwen3embedding], trained using an innovative multi-stage pipeline that combines large-scale unsupervised pretraining with supervised fine-tuning on high quality datasets, demonstrates strong performance across both general purpose and domain-specific retrieval tasks. These advancements highlight the growing effectiveness and specialization of LLM-based embedding models in both general and domain-specific retrieval scenarios, particularly within the legal domain.
3 Preliminary
In the following, a bold lowercase letter denotes a vector, a bold uppercase letter denotes a matrix, a lowercase letter denotes a scalar or a sequence of words, and a scripted uppercase letter denotes a set.
3.1 Task Definition
In legal case retrieval, given the query case $q$, and the set of $n$ candidate cases $\mathcal{D}=\{d_{1},d_{2},...,d_{n}\}$, our task is to retrieve a set of relevant cases $\mathcal{D}^{*}=\{d^{*}_{i}|d^{*}_{i}\in D\wedge relevant(d^{*}_{i},q)\}$, where $relevant(d^{*}_{i},q)$ denotes that $d^{*}_{i}$ is a relevant case of the query case $q$. The relevant cases are called precedents in legal domain, which refer to the historical cases that can support the judgement of the query case.
3.2 Case Graph Construction
In CaseGNN, legal cases are transformed into case graphs to capture their underlying structural information. Specifically, legal relations among entities within a case are extracted using information extraction tools, with entities represented as nodes and the extracted relations as edges in the resulting case graph.
3.2.1 Extraction of Legal Relation and entities.
[…]
3.2.2 Case Graph Construction.
In legal domain, the relevance of cases is primarily determined by the alignment of legal facts and legal issues[promptcase]. Legal facts describe fundamental aspects of a case such as who, when, what, where, and why. And legal issues pertain to the legal disputes between parties that require judicial resolution[promptcase]. Therefore, the case graph of every case is constructed into a legal fact graph and a legal issue graph separately. Firstly, the legal fact and legal issue are extracted by PromptCase module, which is detailed in the PromptCase paper [promptcase]. Given a legal case $c$, let $R_{c,\text{fact}}$ and $R_{c,\text{issue}}$ denote the sets of relation triplets corresponding to the legal fact and legal issue, respectively. Based on these triplets, two text-attributed case graphs are constructed as the legal fact graph $G_{c,\text{fact}}=(V_{c,\text{fact}},E_{c,\text{fact}})$ and the legal issue graph $G_{c,\text{issue}}=(V_{c,\text{issue}},E_{c,\text{issue}})$. In $G_{c,\text{fact}}$, the node set $V_{c,\text{fact}}$ consists of all head and tail entities $h$ and $t$ in $R_{c,\text{fact}}$ represented as nodes $v_{h}$ and $v_{t}$. The edge set $E_{c,\text{fact}}$ includes directed edges $e_{v_{h}v_{t}}$ corresponding to relations $r$ from each head entity $h$ to its associated tail entity $t$. The construction of $G_{c,\text{issue}}$ follows the same procedure using $R_{c,\text{issue}}$. in $R_{c,\text{fact}}$ as $e_{v_{h}v_{t}}$. The same construction process is applied to the $G_{c,\text{issue}}$ for the legal issue. Additionally, each case’s fact graph and issue graph include a virtual global node $v_{\text{g, fact}}$ and $v_{\text{g, issue}}$ representing overall textual semantics, connected to all nodes to facilitate global information propagation.
4 Method
In this section, we present the LEXA framework in detail. The method begins by fine-tuning a retrieval oriented LLM embedding model for informative case representation generation, which is described in Section 4.1. With the proposed LEXA model, the node features are initialised through legal-domain prompts that guide LEXA to encode entities, relations, facts, and issues into text embeddings with rich semantic information, as described in Section 4.2. A Legal Edge Updated Graph Neural Network (LEUGNN), proposed in Section 4.3, refines these embeddings through message passing over legal relations within the case graph. To enhance generalization, graph augmentation is leveraged to construct diverse positive and negative samples for contrastive learning, as described in Section 4.4. Finally, a contrastive learning objective, detailed in Section 4.5, aligns embeddings of augmented views while preserving discriminative information, resulting in robust and semantically rich legal case representations.
Figure 1: The framework of LEXA. Given legal case $c$, the legal fact and the legal issue sections are converted into case graph. The case graph is processed by L layers of EUGAT and a Readout function to obtain an overall case graph representation. The whole framework is trained with the contrastive loss with positive, negative and augmented samples.
4.1 LEXA-8B
To develop a legal retrieval oriented embedding model capable of encoding legal cases into informative representations, the Qwen3-Embedding-8B [qwen3embedding] model was fine-tuned using supervised contrastive learning. The resulting model, LEXA-8B, addresses the base model’s limitation in capturing domain-specific semantics within the legal domain. Each case is first transformed into a structured text input integrating legal facts and legal issues. Through this fine-tuning process, LEXA-8B learns to align embeddings of relevant cases while distinguishing irrelevant ones, enabling it to capture fine-grained semantic relationships and generalise effectively across diverse legal scenarios, thus achieving precise and robust precedent retrieval.
4.1.1 Contextualised Case Encoding
During training, each case is reformulated as $c_{\text{Cont}}$ by concatenating its extracted legal facts and legal issues:
Prompt Template for Contextualised Case Encoding # System Prompt
    The following contains key components of a legal case.
    # User Prompt
    Legal facts: $\{c_{\text{Fact}}\}$.
    Legal issues: $\{c_{\text{Issue}}\}$.
Feeding the reformulated case $c_{\text{Cont}}$ into LEXA-8B, the case is encoded into a dense vector representation as:
$\mathbf{x}_{c}=\text{{LEXA-8B}}(c_{\text{Cont}}),$ (1)
where $\mathbf{x}_{c}\in\mathbb{R}^{d}$ denotes the $d$-dimensional semantic embedding of the case, capturing its key legal element information for downstream retrieval task.
4.1.2 LEXA-8B Fine-tuning Objective
The supervised LLM Fine-tuning objective encourages embeddings of the query case and positive cases to be pulled closer together while pushing embeddings of negative cases apart, thereby enabling the model to learn discriminative and semantically enriched case representations. To further enhance training effectiveness, hard negative samples are introduced by retrieving cases with high BM25 relevance scores that are not labelled as ground-truth matches, ensuring the model is challenged with difficult but informative distinctions:
$\ell=-\text{log}\frac{e^{(s(\mathbf{x}_{q},\mathbf{x}_{d^{+}}))/\tau}}{e^{(s(\mathbf{x}_{q},\mathbf{x}_{d^{+}}))/\tau}+\sum\limits_{i=1}^{n}e^{(s(\mathbf{x}_{q},\mathbf{x}_{d^{easy-}_{i}}))/\tau}+\sum\limits^{m}_{j=1}e^{(s(\mathbf{x}_{q},\mathbf{x}_{d^{hard-}_{j}}))/\tau}}.$ (2)
In Equation 2, $q$ denotes the query case. The positive sample $d^{+}$ is derived from the ground-truth labels. Negative samples are composed of two types: easy negatives $d^{easy-}$, which are randomly drawn from the candidate pool and include in-batch samples; and hard negatives $d^{hard-}$, which are retrieved using high BM25 similarity scores. The variables $n$ and $m$ denote the number of easy and hard negative samples, respectively. The similarity function $s$, implemented using either the dot product or cosine similarity, measures the similarities between case representations. Finally, the temperature coefficient $\tau$ scales the similarity scores, controlling the sharpness of the softmax distribution during contrastive learning.
4.2 Prompt-based Graph Features Initialization with Large Language Models
To generate graph features that capture richer contextual information from legal cases, node and edge features are encoded using large language models (LLMs) to enhance case representation and improving model training. Specifically, prompt is used in text encoding for guiding LLMs to focus on the legal task.
4.2.1 Prompt Templates.
Recent research on prompt-based methods for legal case retrieval [promptcase] demonstrates that incorporating prompts can guide the model to capture richer semantic information, thereby improving performance on downstream tasks. Motivated by this, prompt-based encoding is adopted in feature extraction process. Given an example of relation triplet, “The applicant”, “is”, “a Canadian” and the legal facts and legal issues of the case, the prompts are shown in Table 1. A detail experiment of effectiveness of different prompt templates is described in Section 5.6.
Table 1: Prompt templates for relation triplets features encoding.
Type Text Prompt
Relation Triplets
$\text{prompt}_{(\text{head})}$ “The applicant” “Given the following sentence, The applicant is a Canadian,
the head entity of legal relation triplet is: The applicant”
$\text{prompt}_{(\text{tail})}$ “a Canadian” “Given the following sentence,The applicant is a Canadian,
the tail entity of legal relation triplet is: a Canadian”
$\text{prompt}_{(\text{relation}})$ “is” “Given the following sentence, The applicant is a Canadian,
the relation between two entities of legal relation triplet is: is”
Global Nodes
$\text{prompt}_{(\text{fact})}$ Legal facts text “Legal facts: ” + Legal facts text
$\text{prompt}_{(\text{issue})}$ Legal issues text “Legal issues: ” + Legal issues text
4.2.2 Node and Edge Features of Relation Triplet.
For a head entity node $v_{head}$, tail entity node $v_{tail}$, and relation edge between head and tail entity edge $e_{relation}$ in the case graph, the text encoding is conducted as follows:
$\mathbf{x}_{v_{head}}$ $=\text{LLM}(\text{prompt}_{(\text{head})});$ (3)
$\mathbf{x}_{v_{tail}}$ $=\text{LLM}(\text{prompt}_{(\text{tail})});$ (4)
$\mathbf{x}_{e_{relation}}$ $=\text{LLM}(\text{prompt}_{(\text{relation})}),$ (5)
where $\text{prompt}_{(\text{head})}$, $\text{prompt}_{(\text{tail})}$, and $\text{prompt}_{(\text{relation})}$ are the encoding prompt of node $v_{head}$, node $v_{tail}$ and edge $e_{relation}$ respectively. LLM is a pre-trained text embedding large language model, such as text-embedding-3-large  [^3] https://platform.openai.com/docs/guides/embeddings, E5-Mistral-7B-Instruct [e5mistral] or Qwen3-Embedding-8B [qwen3embedding]. The vectors $\textbf{x}_{v_{head}}\in\mathbb{R}^{d}$, $\textbf{x}_{v_{tail}}\in\mathbb{R}^{d}$, and $\textbf{x}_{e_{relation}}\in\mathbb{R}^{d}$ represent the output of LLM encoding and serve as the feature representations in case graph for node $v_{head}$, $v_{tail}$ and edge $e_{relation}$.
4.2.3 Node and Edge Features of Virtual Global Node.
For node features of the virtual global nodes $v_{\text{g,fact}}$ and $v_{\text{g,issue}}$, the whole texts of legal fact and legal issue are also encoded by the same LLM with prompt. To simplify the feature extraction process, the edge features connecting the global node to other nodes such as node $v$ are directly reused from the feature representation of node $v$:
$\mathbf{x}_{v_{\text{g,fact}}}$ $=\text{LLM}(\text{prompt}_{(\text{fact})});$ (6)
$\mathbf{x}_{e_{\text{g,fact}}}$ $=\mathbf{x}_{v};$ (7)
$\mathbf{x}_{v_{\text{g,issue}}}$ $=\text{LLM}(\text{prompt}_{(\text{issue})});$ (8)
$\mathbf{x}_{e_{\text{g,issue}}}$ $=\mathbf{x}_{v}.$ (9)
4.3 Legal Edge Updated Graph Neural Network
Considering the importance of edges of case graph that represents the pivotal relations between legal entities, a novel Legal Edge Updated Graph Neural Network (LEUGNN) is proposed to incorporate the renewable nodes and edges features simultaneously. The LEUGNN includes two parts, the Edge-updated Graph Attention Layer (EUGAT) and readout function.
4.3.1 Edge-updated Graph Attention Layer.
Edge-updated graph attention layer (EUGAT) is a fundamental component of LEUGNN for transferring the case graph into informative case representation, which includes different node feature update and edge feature update schemes respectively.
[…]
$\mathbf{h}^{{}^{\prime}}_{e_{uv}}=\mathbf{h}_{e_{uv}}+\mathop{\text{Avg}}_{k=1:K}(\text{LeakyReLU}({\mathbf{w}_{\text{att,e}}^{k}}^{T}[\mathbf{W}^{k}_{n}\cdot\mathbf{h}_{v}\mathbin{\|}\mathbf{W}^{k}_{e}\cdot\mathbf{h}_{e_{uv}}\mathbin{\|}\mathbf{W}^{k}_{n}\cdot\mathbf{h}_{u}])),$ (10)
where $\mathbf{h}^{{}^{\prime}}_{e_{uv}}\in\mathbb{R}^{d^{\prime}}$ is the updated edge feature, and Avg means the average of the output vectors of K heads. All the weight matrices W are in $\mathbb{R}^{d^{\prime}\times d}$. Specifically, $\textbf{W}_{n}$ is the neighbour node update weight matrix and $\textbf{W}_{e}$ is the edge update weight matrix that are the same matrices used in Equation 11 and Equation 12, $\mathbf{w}_{\text{att,e}}^{k}\in\mathbb{R}^{3d^{\prime}}$ is the edge weight vector of attention layer. Moreover, to avoid over-smoothing problem, a residual connection is added.
$\bullet$ Node Feature Update. For node feature update, a self-attention module will be used to aggregate the nodes and its neighbour nodes and edges information to an informative representation. Also, a residual connection is added for avoiding over-smoothing. As previous study shows, multi-head attention has better performance than original attention[transformer]. According to multi-head attention mechanism, the update of node $v$ with $K$ attention heads is defined as:
$\mathbf{h}^{{}^{\prime}}_{v}=\mathbf{h}_{v}+\mathop{\text{Avg}}_{k=1:K}(\sum_{u\in N(v)}\alpha^{k}(\mathbf{W}^{k}_{n}\cdot\mathbf{h}_{u}+\mathbf{W}^{k}_{e}\cdot\mathbf{h}_{e_{uv}})),$ (11)
where $\mathbf{h}^{{}^{\prime}}_{v}\in\mathbb{R}^{d^{\prime}}$ is the updated node feature, and Avg means the average of the output vectors of K heads. Specially, the input of the first EUGAT layer is initialised as $\textbf{h}_{v}=\textbf{x}_{v}$, $\textbf{h}_{u}=\textbf{x}_{u}$ and $\textbf{h}_{e_{uv}}=\textbf{x}_{e_{uv}}$. All the weight matrices W are in $\mathbb{R}^{d^{\prime}\times d}$. Specifically, $\textbf{W}_{n}$ is the neighbour node update weight matrix and $\textbf{W}_{e}$ is the edge update weight matrix respectively. $\alpha^{k}$ is the attention weight in the attention layer as:
$\alpha^{k}=\text{Softmax}(\text{LeakyReLU}({\mathbf{w}_{\text{att,n}}^{k}}^{T}[\mathbf{W}^{k}_{n}\cdot\mathbf{h}_{v}\mathbin{\|}\mathbf{W}^{k}_{n}\cdot\mathbf{h}_{u}\mathbin{\|}\mathbf{W}^{k}_{e}\cdot\mathbf{h}_{e_{uv}}])),$ (12)
where Softmax is the softmax function, LeakyReLU is the non-linear function, $\mathbf{w}_{\text{att,n}}^{k}\in\mathbb{R}^{3d^{\prime}}$ is the node weight vector of attention layer and $\mathbin{\|}$ denotes concatenation of vectors.
4.3.2 Readout Function for Legal Case Representation
To generate the graph representation with the updated node and edge features, the Readout function is utilised in many GNN works [GIN, casegnn], which is always defined as:
$\mathbf{h}_{G}$ $=\text{Readout}(h_{v}^{\text{(}final)}|v\in G),$ (13)
where $h_{v}^{\text{(}final)}\in\mathbb{R}^{d^{\prime}}$ denotes the representation of node $v$ in graph $G$ and Readout is an aggregation function that produces a graph-level representation through operations such as max pooling, sum pooling, or attention pooling.
Specifically, in this paper, since the virtual global is connected to all other nodes in the graph, the final iteration feature of the virtual global node $\mathbf{h}_{v_{g}}\in\mathbb{R}^{d^{\prime}}$ can be regarded as the final representation of the case graph due to its full graph connectivity. Accordingly, in the design of LEUGNN, the Readout function in Equation 13 is implemented by directly using the representation of the global nodes $\mathbf{h}_{v_{g,\text{fact}}}\in\mathbb{R}^{d^{\prime}}$ and $\mathbf{h}_{v_{g,\text{issue}}}\in\mathbb{R}^{d^{\prime}}$ respectively:
$\mathbf{h}_{c,\text{fact}}$ $=\mathbf{h}_{v_{g,\text{fact}}},$ (14)
$\mathbf{h}_{c,\text{issue}}$ $=\mathbf{h}_{v_{g,\text{issue}}}.$ (15)
The final representation of case $c$, $\textbf{h}_{c}\in\mathbb{R}^{2d^{\prime}}$ is the concatenation of legal fact graph representation $\textbf{h}_{c,\text{fact}}$ and legal issue graph representation $\textbf{h}_{c,\text{issue}}$, which is defined as:
$\textbf{h}_{c}=\textbf{h}_{c,\text{fact}}\mathbin{\|}\textbf{h}_{c,\text{issue}}.$ (16)
4.4 Graph Augmentation
Graph augmentation is utilised in this paper to generate more positive and negative samples for graph contrastive learning. To effectively obtain the augmented graph while reserving the informative legal structure information, the graph augmentation methods of edge dropping and feature masking are leveraged for graph contrastive learning during the training process.
4.4.1 Edge Dropping
Edge dropping is a popular graph augmentation method that randomly drop the edges of the original graph as a new graph. Given a case graph $G_{c}$, the edge dropping augmentation graph is denoted as $G_{c,\text{drop}}$, where the adjacency matrices of $G_{c}$ and $G_{c,\text{drop}}$ is defined as:
$\mathbf{A}_{c,\text{drop}}=\mathbf{M}\odot\mathbf{A}_{c},$ (17)
where $\textbf{A}_{c}\in\mathbb{R}^{n\times n}$ is the adjacency matrix of the graph $G_{c}$. $\textbf{M}\in{\{0,1\}}^{n\times n}$ is a binary mask matrix that acts on the adjacency matrix for edge dropping, which generates the adjacency matrix $\mathbf{A}_{c,\text{drop}}\in\mathbb{R}^{n\times n}$ for augmented graph. Specifically, $\textbf{M}_{i,j}=\text{Bernoulli}(\epsilon)$, where $\epsilon\in(0,1)$ refers to the drop rate hyperparameter and $\odot$ indicates the Hadamard product (also known as element-wise product).
4.4.2 Feature Masking
Another popular graph augmentation method called feature masking is also leveraged for training in this paper. Specifically, to mask the node and edge features in a graph, the column of node and edge feature tensors will be masked in a probability $p_{\text{node}}\in(0,1)$ and $p_{\text{edge}}\in(0,1)$ respectively, which are two hyperparameters for training.
4.5 Graph Augmentation Contrastive Learning Objective Function
Contrastive learning is a tool that aims at pulling the positive samples closer while pushing the negative samples far away to provide the training signal for training [SAILER, MVCL, CL4LJP]. The goal of training the LEXA model is to distinguish the relevant cases from irrelevant cases given a large case pool, which corresponds to the goal of contrastive learning. In contrast to CaseGNN, LEXA incorporates augmented graphs into the contrastive learning objective to enrich the training signals and enhance representation learning. Accordingly, the objective function for model training is defined as follows:
$\ell=-\text{log}$ (18)
$\left(\frac{e^{\frac{s(\mathbf{h}_{q},\mathbf{h}_{d^{+}})}{\tau}}+e^{\frac{s(\mathbf{h}_{q},\mathbf{h}_{d_{aug}^{+}})}{\tau}}}{e^{\frac{s(\mathbf{h}_{q},\mathbf{h}_{d^{+}})}{\tau}}+e^{\frac{s(\mathbf{h}_{q},\mathbf{h}_{d_{aug}^{+}})}{\tau}}+\sum\limits_{i=1}^{n}e^{\frac{s(\mathbf{h}_{q},\mathbf{h}_{d^{easy-}_{i}})}{\tau}}+\sum\limits^{n}_{i=1}e^{\frac{s(\mathbf{h}_{q},\mathbf{h}_{d^{easy-}_{i,aug}})}{\tau}}+\sum\limits^{m}_{j=1}e^{\frac{s(\mathbf{h}_{q},\mathbf{h}_{d^{hard-}_{j}})}{\tau}}}\right).$
In equation 18, $q$ denotes the query case. The positive sample $d^{+}$ is directly derived from the ground truth labels in the dataset. Easy negative samples $d^{easy-}$ are drawn randomly from the entire candidate pool and also include in-batch samples from other queries. To further enhance training effectiveness, hard negative samples $d^{hard-}$ are incorporated by identifying challenging cases that are difficult to distinguish from relevant ones. A candidate case is considered a hard negative if it receives a high BM25 relevance score but is not labelled as a ground truth match. Such cases exhibit high textual similarity to the query while remaining irrelevant, thereby offering more informative contrastive signals. The variables $n$ and $m$ represent the number of easy and hard negative samples, respectively. The similarity function $s$, which measures the closeness between representations, can be implemented using either the dot product or cosine similarity. The temperature coefficient $\tau$ is used to scale the similarity scores, controlling the sharpness of the softmax distribution during contrastive learning.
Specifically, in the LEXA model, graph augmentation is applied to both positive and negative cases to provide additional training signals, addressing the challenges posed by limited and costly legal data annotation. In Equation 18, the augmented graph of relevant cases $d^{+}$ is denoted as $d_{aug}^{+}$ while the augmented graph of easy negative sample $d^{easy-}$ is denoted as $d_{aug}^{easy-}$. During the graph contrastive learning process, only positive and easy negative samples are augmented; hard negative samples $d^{hard-}$ are used in their original form without augmentation. This design is motivated by experimental observations indicating that a limited number of hard negative samples is sufficient for effective training, whereas an excessive number can lead to notable performance degradation. Furthermore, for each case graph, only a single augmentation strategy is employed to generate an augmented graph. This decision is motivated by the empirical finding that using multiple augmentation methods simultaneously results in reduced performance. Additionally, during training, either augmented positive samples or augmented easy negative samples are used individually, as combining both was found to yield suboptimal results. The complete model pipeline is illustrated in Fig. 1.
5 Experiments
In this section, the following research questions (RQs) will be studied extensively to evaluate the effectiveness of the proposed LEXA model:
  • RQ1: How does LEXA perform compared with other legal case retrieval models?
  • RQ2: How effective are edge feature update and graph augmentation in LEXA?
  • RQ3: How effective is EUGAT compared with other GNNs?
  • RQ4: How effective is prompt-based LLM features encoding?
  • RQ5: How does graph contrastive learning help with LCR in LEXA?
  • RQ6: How do hyper-parameter settings affect LEXA?
5.1 Setup
In setup, the detailed descriptions of two benchmark datasets, seven evaluated metrics, seven LCR baselines and the experiment implementations are provided.
5.1.1 Datasets.
To assess the effectiveness of the proposed LEXA model, experiments are conducted on two benchmark legal case retrieval datasets, COLIEE2022 [COLIEE2022] and COLIEE2023 [COLIEE2023], both released as part of the Competition on Legal Information Extraction and Entailment (COLIEE). These datasets consist of legal cases sourced from the Federal Court of Canada. The task of legal case retrieval is retrieving relevant cases from the full candidate pool given a query case. Although the training sets of COLIEE2022 and COLIEE2023 partially overlap, their test sets are entirely distinct. Additionally, as shown in Table 2, the average number of relevant cases per query differs between the two datasets, resulting in varying levels of retrieval difficulty. These datasets represent the most widely used English benchmarks for legal case retrieval. Notably, LEXA can be readily adapted to multilingual settings through appropriate legal information extraction tools and language models.
Table 2: Statistics of datasets.
Datasets COLIEE2022 COLIEE2023
train test train test
# Query 898 300 959 319
# Candidates 4415 1563 4400 1335
# Avg. relevant cases 4.68 4.21 4.68 2.69
Avg. length (# token) 6724 6785 6532 5566
Largest length (# token) 127934 85136 127934 61965
5.1.2 Baselines.
According to the recent research [promptcase, SAILER], six popular and state-of-the-art methods are compared as well as the competition winners:
  • BM25 [BM25] is a classic term-matching model that remains a strong competitor in many retrieval tasks. Building on the TF–IDF framework [TF-IDF], BM25 incorporates document length normalization and a corpus-level length prior to mitigate biases arising from varying document sizes.
  • LEGAL-BERT [LEGAL-BERT] is a bert-based model pre-trained on a large and diverse corpus of English legal text spanning legislation and judicial opinions from multiple jurisdictions. Its domain-specific pre-training makes it a strong benchmark for legal NLP tasks.
  • MonoT5 [monot5] adapts the T5 text-to-text architecture [T5] to ranking scenarios. By framing relevance estimation as a generation problem, MonoT5 demonstrates powerful language modelling capabilities in retrieval settings.
  • SAILER [SAILER] is a structure-aware legal language model designed specifically for legal case retrieval. Its encoder–decoder architecture is trained to capture the internal structure of cases, enabling it to achieve competitive retrieval performance.
  • PromptCase [promptcase] is a prompt-based reformulation method that guides language models to focus on key legal aspects. By explicitly encoding legal facts and issues, it enhances the model’s ability to retrieve cases that match the key legal elements of a query.
  • E5-Mistral-7B-Instruct [e5mistral]is a general purpose embedding model built through instruction tuning on a broad set of corpora. It produces strong semantic embeddings and is widely used for retrieval and semantic search tasks.
  • Qwen3-Embedding-8B [qwen3embedding] generates high-quality representations suitable for similarity and retrieval applications. Its extensive training and robust generalization make it a reliable baseline even in specialized domains such as law.
  • Inf-Retriever-V1 [inf-retriever-v1] is optimized for large-scale retrieval workloads, focusing on creating highly discriminative embeddings that balance retrieval accuracy with computational efficiency.
  • CaseGNN [casegnn] models each case as a graph enriched with textual attributes, using a graph neural network and contrastive learning to derive case representations. It serves as an important reference point for graph-based methods for legal case retrieval task.
  • CaseLink [caselink] advances legal case retrieval by incorporating connections among cases. By leveraging inter-case links, it captures latent legal dependencies that further enhance retrieval quality.
5.1.3 Metrics.
In this experiment, seven metrics that often utilised for evaluating information retrieval models and LCR models are leveraged to comprehensively evaluate the LEXA model. According to the previous LCR works [promptcase, LeCaRD, SAILER], top 5 ranking results are evaluated and all metrics are the higher the better. The details of the seven metrics are as follows:
  • Precision (P) captures how reliably a model identifies relevant items. It is defined as: $\text{Precision}={\text{TP}}/({\text{TP}+\text{FP}})$, where TP represents correctly retrieved relevant cases, and FP counts the cases incorrectly retrieved as relevant.
  • Recall (R) evaluates the extent to which a model successfully retrieves all relevant items. It is calculated by $\text{Recall}=\text{TP}/(\text{TP}+\text{FN})$, with FN denoting relevant cases the model failed to retrieve.
  • Micro F1 (Mi-F1) offers a unified measure that balances precision and recall by computing their harmonic mean across all instances: $\text{Mi-F1}=2\times(\text{Precision}\times\text{Recall})/(\text{Precision}+\text{Recall}))$.
  • Macro F1 (Ma-F1) assesses performance in a more class-agnostic manner by averaging the F1 score of each class, ensuring equal contribution regardless of class size: $\text{Ma-F1}=\frac{1}{N}\sum_{i=1}^{N}\text{F1}_{i}$.
  • Mean Reciprocal Rank (MRR) @K measures how quickly a system retrieves the first relevant item. It is defined as $\text{MRR}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{\text{rank}_{i}}$, where $N$ is the number of queries, $\text{rank}_{i}$ is the rank position of the first correct result for the $i$-th query.
  • Mean Average Precision (MAP) @K captures the average retrieval precision across all queries. For each query, average precision is computed separately, and then aggregated as $\text{MAP}=\frac{1}{N}\sum_{i=1}^{N}\text{AP}_{i}$.
  • Normalized Discounted Cumulative Gain (NDCG) @K examines ranking quality by rewarding relevant items appearing near the top while penalizing those placed lower. This measure is computed as $\text{NDCG@K}=\text{DCG@K}/\text{IDCG@K}$, where DCG captures the discounted gain from the actual ranked list, and IDCG represents the best possible ordering within the top K positions.
5.1.4 Implementation.
The implementation includes data preprocessing and hyperparameters.
[…]
Table 3: Overall performance on COLIEE2022 (%). Underlined numbers indicate the best baselines. Bold numbers indicate the best performance of all methods. Both one-stage and two-stage Top-5 results are reported.
Methods COLIEE2022
P@5 R@5 Mi-F1 Ma-F1 MRR@5 MAP NDCG@5
One-stage
BM25 17.9 21.2 19.4 21.4 23.6 25.4 33.6
LM
LEGAL-BERT 4.47 5.30 4.85 5.38 7.42 7.47 10.9
MonoT5 0.71 0.65 0.60 0.79 1.39 1.41 1.73
SAILER 16.6 15.2 14.0 16.8 17.2 18.5 25.1
PromptCase 17.1 20.3 18.5 20.5 35.1 33.9 38.7
LLM
E5-Mistral-7B-Instruct 21.4 25.4 23.2 25.7 26.8 28.5 38.0
Qwen3-Embedding-8B 21.6 25.7 23.5 26.0 26.7 29.0 37.8
Inf-Retriever-V1 21.1 25.1 22.9 25.5 26.6 28.7 37.9
Graph
CaseGNN 35.5 $\pm$0.2 42.1 $\pm$0.2 38.4 $\pm$0.3 42.4 $\pm$0.1 66.8 $\pm$0.8 64.4 $\pm$0.9 69.3 $\pm$0.8
CaseLink 37.0 $\pm$0.1 43.9 $\pm$0.1 40.1 $\pm$0.1 44.2 $\pm$0.1 67.3 $\pm$0.5 65.0 $\pm$0.2 70.3 $\pm$0.1
LEXA (Ours) 43.7 $\pm$0.2 51.9 $\pm$0.2 47.5 $\pm$0.2 52.4 $\pm$0.3 77.0 $\pm$0.5 74.7 $\pm$0.9 79.3 $\pm$0.8
Two-stage
LM
SAILER 23.8 25.7 24.7 25.2 43.9 42.7 48.4
PromptCase 23.5 25.3 24.4 30.3 41.2 39.6 45.1
LLM
E5-Mistral-7B-Instruct 22.0 26.2 24.0 25.8 48.4 46.8 51.5
Qwen3-Embedding-8B 21.9 26.0 23.8 25.8 48.8 48.0 52.1
Inf-Retriever-V1 22.2 26.4 24.1 26.1 45.5 44.5 49.6
Graph
CaseGNN 22.9 $\pm$0.1 27.2 $\pm$0.1 24.9 $\pm$0.1 27.0 $\pm$0.1 54.9 $\pm$0.4 54.0 $\pm$0.5 57.3 $\pm$0.6
CaseLink 24.7 $\pm$0.1 29.1 $\pm$0.1 26.8 $\pm$0.1 29.2 $\pm$0.1 56.0 $\pm$0.2 55.0 $\pm$0.2 58.6 $\pm$0.1
LEXA (Ours) 25.4 $\pm$0.2 30.2 $\pm$0.2 27.6 $\pm$0.2 30.0 $\pm$0.1 59.0 $\pm$0.6 58.0 $\pm$0.8 61.2 $\pm$0.6
Table 4: Overall performance on COLIEE2023 (%). Underlined numbers indicate the best baselines. Bold numbers indicate the best performance of all methods. Both one-stage and two-stage Top-5 results are reported.
Methods COLIEE2023
P@5 R@5 Mi-F1 Ma-F1 MRR@5 MAP NDCG@5
One-stage
BM25 16.5 30.6 21.4 22.2 23.1 20.4 23.7
LM
LEGAL-BERT 4.64 8.61 6.03 6.03 11.4 11.3 13.6
MonoT5 0.38 0.70 0.49 0.47 1.17 1.33 0.61
SAILER 12.8 23.7 16.6 17.0 25.9 25.3 29.3
PromptCase 16.0 29.7 20.8 21.5 32.7 32.0 36.2
LLM
E5-Mistral-7b-Instruct 16.0 29.7 20.8 21.5 21.9 22.9 30.8
Qwen3-Embedding-8B 18.4 34.1 23.9 25.2 25.8 27.5 36.4
Inf-Retriever-V1 19.0 35.3 24.7 26.0 26.5 28.0 37.6
Graph
CaseGNN 17.7 $\pm$0.7 32.8 $\pm$0.7 23.0 $\pm$0.5 23.6 $\pm$0.5 38.9 $\pm$1.1 37.7 $\pm$0.8 42.8 $\pm$0.7
CaseLink 20.9 $\pm$0.3 38.4 $\pm$0.6 27.1 $\pm$0.3 28.2 $\pm$0.3 45.8 $\pm$0.5 44.3 $\pm$0.7 49.8 $\pm$0.4
LEXA 22.6 $\pm$0.4 41.9 $\pm$0.8 29.3 $\pm$0.6 30.5 $\pm$0.7 48.7 $\pm$0.7 47.0 $\pm$0.7 52.6 $\pm$0.8
Two-stage
LM
SAILER 19.6 32.6 24.5 23.5 37.3 36.1 40.8
PromptCase 21.8 36.3 27.2 26.5 39.9 38.7 44.0
LLM
E5-Mistral-7b-Instruct 19.7 36.6 25.6 26.5 46.9 44.4 49.9
Qwen3-Embedding-8B 20.1 37.4 26.2 27.0 47.7 46.3 51.1
Inf-Retriever-V1 19.7 36.6 25.6 26.6 48.2 46.9 51.6
Graph
CaseGNN 20.2 $\pm$0.2 37.6 $\pm$0.5 26.3 $\pm$0.3 27.3 $\pm$0.2 45.8 $\pm$0.9 44.4 $\pm$0.8 49.6 $\pm$0.8
CaseLink 21.0 $\pm$0.3 38.9 $\pm$0.5 27.1 $\pm$0.3 28.2 $\pm$0.3 48.8 $\pm$0.2 47.2 $\pm$0.1 52.6 $\pm$0.1
LEXA 21.3 $\pm$0.2 39.5 $\pm$0.3 27.7 $\pm$0.2 28.6 $\pm$0.2 50.6 $\pm$0.2 49.3 $\pm$0.2 54.1 $\pm$0.2
5.2 Overall Performance (RQ1)
The results in Table 3 and Table 4 reveal clear performance patterns across both datasets. Traditional retrieval methods and language model baselines occupy the lower end of the accuracy performance, while methods that incorporate structural information especially graph-based models achieve substantially better retrieval quality. Across every evaluation setting, LEXA delivers the strongest results and consistently establishes new state of the art performance.
On COLIEE2022, BM25 and language model based systems such as LEGAL-BERT, MonoT5, SAILER, and PromptCase achieve modest accuracy and weak ranking performance. The gap between these models and large language model embeddings widens considerably. E5-Mistral-7B-Instruct, Qwen3-Embedding-8B, and Inf-Retriever-V1 provide a noticeable boost, surpassing all language model baselines by large margins. However, none of these embedding based methods come close to matching the performance of the graph-based models. CaseGNN and CaseLink demonstrate substantial improvements over language model and large language model methods, underscoring the advantage of explicitly modelling legal entities and relations. LEXA further extends this advantage, outperforming CaseLink by large absolute margins. These gains indicate a significant advancement in retrieval capability.
A similar trend appears on COLIEE2023. The language model and large language model baselines again show limited ability to capture the structural reasoning required for legal case retrieval. In contrast, graph-based models dominate the upper end of the performance distribution. CaseLink continues to be the strongest baseline on this dataset across nearly all metrics. Yet LEXA surpasses CaseLink on every measure by a consistent and meaningful amount, setting new records in both precision oriented and ranking oriented metrics.
The two stage retrieval pipeline further highlights the relative strengths of the different approaches. While language model and large language model methods benefit moderately from using BM25 as a first stage filter, their improvements remain limited. Graph-based approaches, however, show consistent gains especially on COLIEE2023 where the candidate set provided by BM25 is more informative. Even in scenarios where CaseLink narrows the gap, LEXA remains the strongest method across all metrics and both datasets. The improvements observed in the two-stage setting confirm that LEXA is effective not only in an end-to-end retrieval scenario but also when functioning as a re-ranking model.
Overall, the experimental results demonstrate that methods relying solely on lexical similarity or sentence level semantics are insufficient for capturing the complex relational structure of legal cases. Graph-based models provide a clear advantage, and LEXA by combining prompt-guided initialization of node representations with graph message passing and contrastive graph augmentation achieves the highest accuracy, recall, and ranking performance among all evaluated methods.
Table 5: Ablation study. (%)
Methods COLIEE2022
P@5 R@5 Mi-F1 Ma-F1 MRR@5 MAP NDCG@5
PromptCase 17.1 20.3 18.5 20.5 35.1 33.9 38.7
w/o $v_{g}$ 1.6 $\pm$0.1 2.9 $\pm$0.1 2.1 $\pm$0.1 2.2 $\pm$0.1 4 $\pm$0.1 4.0 $\pm$0.1 4.8 $\pm$0.2
Avg Readout 30.5 $\pm$0.5 36.2 $\pm$0.6 33.1 $\pm$0.5 36.8 $\pm$0.5 61.3 $\pm$0.4 59.0 $\pm$0.1 64.6 $\pm$0.5
CaseGNN 35.5 $\pm$0.2 42.1 $\pm$0.2 38.4 $\pm$0.3 42.4 $\pm$0.1 66.8 $\pm$0.8 64.4 $\pm$0.9 69.3 $\pm$0.8
+GCL 35.5 $\pm$0.3 42.2 $\pm$0.4 38.6 $\pm$0.3 42.5 $\pm$0.6 67.1 $\pm$0.7 64.6 $\pm$0.1 70.2 $\pm$0.5
+EUGAT 35.9 $\pm$0.5 42.7 $\pm$0.6 39.2 $\pm$0.7 43.3 $\pm$0.5 67.2 $\pm$0.9 64.0 $\pm$1.2 69.4 $\pm$1.2
w/o Tuned LLM 36.5 $\pm$0.6 43.3 $\pm$0.7 39.6 $\pm$0.6 43.8 $\pm$0.7 68.1 $\pm$1.1 65.3 $\pm$1.1 70.8 $\pm$1.1
LEXA 43.7 $\pm$0.2 51.9 $\pm$0.2 47.5 $\pm$0.2 52.4 $\pm$0.3 77.0 $\pm$0.5 74.7 $\pm$0.9 79.3 $\pm$0.8
Methods COLIEE2023
P@5 R@5 Mi-F1 Ma-F1 MRR@5 MAP NDCG@5
PromptCase 16.0 29.7 20.8 21.5 32.7 32.0 36.2
w/o $v_{g}$ 1.6 $\pm$0.1 2.9 $\pm$0.1 2.1 $\pm$0.1 2.2 $\pm$0.1 4.0 $\pm$0.1 2.9 $\pm$0.1 4.8 $\pm$0.1
Avg Readout 17.6 $\pm$0.4 32.6 $\pm$0.7 22.8 $\pm$0.5 23.6 $\pm$0.4 37.7 $\pm$1.0 36.6 $\pm$0.7 41.7 $\pm$0.5
CaseGNN 17.7 $\pm$0.7 32.8 $\pm$0.7 23.0 $\pm$0.5 23.6 $\pm$0.5 38.9 $\pm$1.1 37.7 $\pm$0.8 42.8 $\pm$0.7
+GCL 17.7 $\pm$0.1 32.9 $\pm$0.2 23.0 $\pm$0.1 23.7 $\pm$0.2 39.1 $\pm$0.3 37.8 $\pm$0.2 42.9 $\pm$0.2
+EUGAT 18.0 $\pm$0.3 33.4 $\pm$0.5 23.4 $\pm$0.3 24.0 $\pm$0.2 39.3 $\pm$0.7 38.1 $\pm$0.7 43.2 $\pm$0.5
w/o Tuned LLM 18.2 $\pm$0.3 33.8 $\pm$0.4 23.7 $\pm$0.4 24.3 $\pm$0.3 40.0 $\pm$0.2 38.9 $\pm$0.3 43.8 $\pm$0.3
LEXA 22.6 $\pm$0.4 41.9 $\pm$0.8 29.3 $\pm$0.6 30.5 $\pm$0.7 48.7 $\pm$0.7 47.0 $\pm$0.7 52.6 $\pm$0.8
5.3 Ablation Study (RQ2)
The ablation analysis in Table 5 evaluates a series of progressively enhanced model variants in order to understand how each design choice contributes to LEXA. (1) PromptCase, a purely text based setting without graph; (2) using the case graph that excludes the virtual global node (w/o $v_{g}$); (3) using the average pooling in CaseGNN (Avg Readout); (4) CaseGNN; (5) extending CaseGNN by incorporating graph contrastive learning (+GCL); (6) changing CaseGNN GNN layer from EdgeGAT to EUGAT (+EUGAT); (7) The full LEXA system combines both enhancements in a unified framework. All variants are tested on the two datasets using every evaluation measure in a single stage retrieval setting.
The ablation results in Table 5 provide a clear picture of how each component contributes to the performance of LEXA. Across both COLIEE2022 and COLIEE2023, the progression from simple text based representations to fully structured graph learning highlights a consistent trend: the richer the modelling of case structure and node interactions, the stronger the retrieval ability. PromptCase offers a text only baseline that relies solely on the encoded fact and issue statements. While reasonably strong for a non-graph method, its performance is far below any configuration that incorporates graph structure. Removing the global node leads to a dramatic collapse in effectiveness on both datasets. The near zero values across all metrics indicate that graph message passing without a meaningful semantic anchor prevents the model from forming any coherent case representation. Introducing even a simple aggregation mechanism substantially shifts the landscape. The Avg Readout variant, which pools node embeddings by averaging, provides a sizeable jump compared to PromptCase on both datasets. This confirms that the structural cues present in the case graphs carry essential information, even when used in a minimal way. Enhancements to CaseGNN further highlight the importance of more expressive graph reasoning. Adding contrastive learning (+GCL) consistently improves ranking metrics such as MRR@5, MAP, and NDCG@5, showing that the additional training objective provides stronger supervision for distinguishing subtle differences between similar cases. Replacing EdgeGAT with EUGAT (+EUGAT) yields another set of gains, reflecting the benefits of more flexible and informative edge updates. The improvements are visible across nearly all metrics in both dataset. The variant that removes the tuned large language model (w/o Tuned LLM) sits between the enhanced CaseGNN variants and the full LEXA model. Although it performs competitively, its results show that high quality node initialization remains important, especially for capturing fine-grained semantic distinctions that downstream graph layers build upon.
At the top of the comparison, LEXA achieves the strongest performance by combining all proposed components, including EUGAT and graph contrastive learning. LEXA outperforms every ablated configuration by large margins. For instance, on COLIEE2022 it surpasses the best non-LEXA variant by more than 9% in NDCG@5, and the improvements on COLIEE2023 are similarly substantial. These consistent gains demonstrate that the combination of expressive graph modelling, semantically grounded global representations, and contrastive training produces a significantly more informative case representation.
Table 6: Effectiveness of GNNs. (%)
Methods COLIEE2022
P@5 R@5 Mi-F1 Ma-F1 MRR@5 MAP NDCG@5
GCN 21.3 $\pm$0.3 25.3 $\pm$0.4 23.2 $\pm$0.2 26.0 $\pm$0.4 46.0 $\pm$0.2 44.4 $\pm$0.1 49.7 $\pm$0.3
GAT 29.3 $\pm$0.1 34.8 $\pm$0.3 31.8 $\pm$0.1 35.3 $\pm$0.2 59.2 $\pm$0.5 56.9 $\pm$0.3 62.3 $\pm$0.7
EdgeGAT 35.5 $\pm$0.2 42.1 $\pm$0.2 38.4 $\pm$0.3 42.4 $\pm$0.1 66.8 $\pm$0.8 64.4 $\pm$0.9 69.3 $\pm$0.8
EUGAT 43.7 $\pm$0.2 51.9 $\pm$0.2 47.5 $\pm$0.2 52.4 $\pm$0.3 77.0 $\pm$0.5 74.7 $\pm$0.9 79.3 $\pm$0.8
Methods COLIEE2023
P@5 R@5 Mi-F1 Ma-F1 MRR@5 MAP NDCG@5
GCN 12.8 $\pm$0.2 23.7 $\pm$0.3 16.5 $\pm$0.1 16.9 $\pm$0.2 27.8 $\pm$0.8 26.8 $\pm$0.6 31.6 $\pm$0.7
GAT 17.4 $\pm$0.3 32.2 $\pm$0.5 22.5 $\pm$0.3 23.1 $\pm$0.5 37.5 $\pm$0.4 36.4 $\pm$0.4 41.4 $\pm$0.4
EdgeGAT 17.7 $\pm$0.7 32.8 $\pm$0.7 23.0 $\pm$0.5 23.6 $\pm$0.5 38.9 $\pm$1.1 37.7 $\pm$0.8 42.8 $\pm$0.7
EUGAT 22.6 $\pm$0.4 41.9 $\pm$0.8 29.3 $\pm$0.6 30.5 $\pm$0.7 48.7 $\pm$0.7 47.0 $\pm$0.7 52.6 $\pm$0.8
5.4 Effectiveness of GNNs (RQ3)
To investigate the contribution of the EUGAT layer, EUGAT is replaced with several commonly used GNN layers of GCN [GCN], GAT [GAT], and EdgeGAT. Results on both datasets across all evaluation metrics are summarized in Table 6.
EUGAT distinguishes itself from previous graph attention mechanisms through its layer-wise edge update strategy. While EdgeGAT incorporates edge features when updating node embeddings, it does not revise those edge features across layers. In contrast, EUGAT continuously updates edge representations, enabling the model to iteratively refine the semantic information carried by each relation. This capability is crucial for legal case graphs, where edges encode the nuanced legal relationships between entities. By progressively enhancing these relation embeddings, EUGAT equips LEXA with more expressive and discriminative edge representations. As a result, the model can better capture the relational patterns that determine case categories, ultimately improving its ability to retrieve the most relevant cases for a given query.
5.5 Effectiveness of Different LLM for Feature Encoding (RQ4)
Table 7: Performance of different LLMs used for feature encoding on COLIEE2022 and COLIEE2023 (%).
Methods COLIEE2022
P@5 R@5 Mi-F1 Ma-F1 MRR@5 MAP NDCG@5
E5-Mistral-7B-Instruct 41.8 $\pm$0.7 50.1 $\pm$0.4 45.8 $\pm$0.3 50.6 $\pm$0.5 73.7 $\pm$0.2 71.1 $\pm$0.1 76.5 $\pm$0.2
Qwen3-Embedding-8B 43.2 $\pm$0.3 51.3 $\pm$0.4 46.9 $\pm$0.4 51.6 $\pm$0.3 75.8 $\pm$0.2 73.5 $\pm$0.2 78.0 $\pm$0.2
LEXA 43.7 $\pm$0.2 51.9 $\pm$0.2 47.5 $\pm$0.2 52.4 $\pm$0.3 77.0 $\pm$0.5 74.7 $\pm$0.9 79.3 $\pm$0.8
Methods COLIEE2023
P@5 R@5 Mi-F1 Ma-F1 MRR@5 MAP NDCG@5
E5-Mistral-7B-Instruct 20.9 $\pm$0.1 38.7 $\pm$0.1 27.1 $\pm$0.1 28.1 $\pm$0.1 45.1 $\pm$0.2 43.6 $\pm$0.1 49.4 $\pm$0.1
Qwen3-Embedding-8B 22.2 $\pm$0.1 41.2 $\pm$0.1 28.8 $\pm$0.1 29.9 $\pm$0.1 46.9 $\pm$0.1 45.6 $\pm$0.3 51.2 $\pm$0.1
LEXA 22.6 $\pm$0.4 41.9 $\pm$0.8 29.3 $\pm$0.6 30.5 $\pm$0.7 48.7 $\pm$0.7 47.0 $\pm$0.7 52.6 $\pm$0.8
In this section, the effectiveness of various LLMs for feature encoding is evaluated. Both E5-Mistral-7B-Instruct and Qwen3-Embedding-8B are employed to encode relation triplets into node and edge feature embeddings. Specifically, Qwen3-Embedding-8B is adopted as the feature encoder in LEXA.
On COLIEE2022, the improvements are relatively modest but consistent. For example, Qwen3-Embedding-8B achieves a higher MAP and NDCG@5, reflecting stronger ranking quality. Similar marginal gains are observed in precision, recall, and F1 scores, indicating that Qwen3-Embedding-8B provides more effective embeddings for both node and edge features. On COLIEE2023, the advantage of Qwen3-Embedding-8B becomes more pronounced. It achieves 2.5% improvement in R@5, leading to higher micro-F1 and macro-F1. The improvements in MAP and NDCG@5 also suggest that Qwen3-Embedding-8B yields more reliable retrieval rankings. Given that COLIEE2023 is a more challenging dataset, these consistent gains highlight the robustness and generalization capability of Qwen3-Embedding-8B.
5.6 Effectiveness of Prompt-based LLM Feature Encoding (RQ5)
Table 8: Prompt templates for relation triplets features encoding.
Type Prompt Templates
prompt 0 $\text{prompt}_{(\text{head})}$ Given the following sentence, The applicant is a Canadian,
the head entity of legal relation triplet is: The applicant
$\text{prompt}_{(\text{tail})}$ Given the following sentence, The applicant is a Canadian,
the tail entity of legal relation triplet is: a Canadian
$\text{prompt}_{(\text{relation})}$ Given the following sentence, The applicant is a Canadian,
the relation between two entities of legal relation triplet is: is
$\text{prompt}_{(\text{fact})}$ Legal facts:
$\text{prompt}_{(\text{issue})}$ Legal issues:
prompt 1 $\text{prompt}_{(\text{head})}$ You are a legal knowledge extraction system tasked with constructing
knowledge graphs from legal texts. Please identify and encode the following
head entity that captures the meaningful legal relationships in the case:
$\text{prompt}_{(\text{tail})}$ You are a legal knowledge extraction system tasked with constructing
knowledge graphs from legal texts. Please identify and encode the following
tail entity that captures the meaningful legal relationships in the case:
$\text{prompt}_{(\text{relation})}$ You are a legal knowledge extraction system tasked with constructing
knowledge graphs from legal texts. Please identify and encode the following
relations that captures the meaningful legal relationships in the case:
$\text{prompt}_{(\text{fact})}$ You are a legal assistant trained to extract and encode structured legal
information from case texts, identify and encode the following legal fact:
$\text{prompt}_{(\text{issue})}$ You are a legal assistant trained to extract and encode structured legal
information from case texts, identify and encode the following legal issue:
prompt 2 $\text{prompt}_{(\text{head})}$ Given the following sentence, The applicant is a Canadian,
the head entity of legal relation triplet is:
$\text{prompt}_{(\text{tail})}$ Given the following sentence, The applicant is a Canadian,
the tail entity of legal relation triplet is:
$\text{prompt}_{(\text{relation})}$ Given the following sentence, The applicant is a Canadian,
the relation between two entities of legal relation triplet is
$\text{prompt}_{(\text{fact})}$ Legal facts:
$\text{prompt}_{(\text{issue})}$ Legal issues:
prompt 3 $\text{prompt}_{(\text{head})}$ Instruction: Encode the legal head entity, The applicant,
from the following legal sentence: The applicant is a Canadian.
$\text{prompt}_{(\text{tail})}$ Instruction: Encode the legal tail entity, a Canadian,
from the following legal sentence: The applicant is a Canadian.
$\text{prompt}_{(\text{relation})}$ Instruction: Encode the legal relation, is,
from the following legal sentence: The applicant is a Canadian.
$\text{prompt}_{(\text{fact})}$ Legal facts:
$\text{prompt}_{(\text{issue})}$ Legal issues:
Table 9: Effectiveness of different prompt templates. Details of prompt templates are shown in Table 8
Methods COLIEE2022
P@5 R@5 Mi-F1 Ma-F1 MRR@5 MAP NDCG@5
no prompt 39.8 $\pm$1.1 47.3 $\pm$1.3 43.3 $\pm$1.2 47.9 $\pm$1.4 71.6 $\pm$2.2 69.1 $\pm$2.1 74.7 $\pm$2.0
prompt 0 43.7 $\pm$0.2 51.9 $\pm$0.2 47.5 $\pm$0.2 52.4 $\pm$0.3 77.0 $\pm$0.5 74.7 $\pm$0.9 79.3 $\pm$0.8
prompt 1 41.3 $\pm$0.9 49.0 $\pm$1.0 44.8 $\pm$1.0 49.4 $\pm$1.1 74.0 $\pm$0.6 71.6 $\pm$0.2 76.6 $\pm$0.3
prompt 2 41.8 $\pm$0.6 49.6 $\pm$0.7 45.4 $\pm$0.6 50.1 $\pm$0.5 74.7 $\pm$1.5 71.9 $\pm$1.1 77.2 $\pm$0.8
prompt 3 40.5 $\pm$1.7 48.1 $\pm$2.0 43.9 $\pm$1.8 48.3 $\pm$2.0 73.0 $\pm$2.6 70.2 $\pm$2.3 75.7 $\pm$1.9
Methods COLIEE2023
P@5 R@5 Mi-F1 Ma-F1 MRR@5 MAP NDCG@5
no prompt 22.0 $\pm$0.5 40.8 $\pm$0.8 28.6 $\pm$0.6 29.6 $\pm$0.8 47.1 $\pm$1.1 45.5 $\pm$1.2 51.0 $\pm$1.3
prompt 0 22.6 $\pm$0.4 41.9 $\pm$0.8 29.3 $\pm$0.6 30.5 $\pm$0.7 48.7 $\pm$0.7 47.0 $\pm$0.7 52.6 $\pm$0.8
prompt 1 21.7 $\pm$0.6 40.3 $\pm$1.1 28.2 $\pm$0.8 29.3 $\pm$0.9 48.9 $\pm$0.5 47.3 $\pm$0.5 52.5 $\pm$0.9
prompt 2 21.7 $\pm$0.2 40.3 $\pm$0.4 28.2 $\pm$0.3 29.2 $\pm$0.4 47.4 $\pm$1.5 45.9 $\pm$1.4 51.2 $\pm$1.1
prompt 3 21.8 $\pm$0.5 40.5 $\pm$0.9 28.3 $\pm$0.6 29.4 $\pm$0.7 46.1 $\pm$0.2 47.7 $\pm$0.3 51.6 $\pm$0.3
This section examines the effectiveness of prompt-based LLM feature encoding in legal case retrieval. Firstly, the impact of alternative prompt templates is assessing on the COLIEE2022 and COLIEE2023 datasets. Given an example of relation triplet, ‘The applicant’, ‘is’, ‘a Canadian’ and the legal facts and legal issues of the case, four prompt templates are shown in the Table 8. And the performance of different prompt templates is shown in Table 9.
On COLIEE2022, applying prompts leads to clear and consistent improvements over the no-prompt baseline across all evaluation metrics. Among the four templates, prompt 0 achieves the strongest performance overall. This suggests that a concise, direct, and minimally intrusive prompt is most effective for guiding LLMs to extract case relevant semantic features. Prompts 1 and 2 also improve over the no-prompt setting but fall slightly behind prompt 0, indicating that additional verbosity or narrative structure does not necessarily yield better retrieval representations. Prompt 3 underperforms relative to other templates, showing that not all forms of instructive prompting are equally beneficial for this task.
While on COLIEE2023, the impact of prompting is notably smaller. All prompt templates perform within a narrow margin. This suggests that, for this dataset, the LLM is less sensitive to prompt wording, likely due to dataset characteristics such as shorter cases or more straightforward relation structures that reduce reliance on prompt-induced guidance. Even so, prompt-based variants remain comparable to or slightly better than the baseline, demonstrating that prompts do not harm performance even when their benefits are limited. These results show the sensitivity of prompt templates and the importance of selecting concise and well structured templates.
5.7 Effectiveness of Graph Augmentation for Graph Contrastive Learning (RQ6)
This experiment evaluates the effectiveness of graph contrastive learning (GCL) by comparing different graph augmentation strategies including node feature masking, edge feature masking, and edge dropping on both COLIEE2022 and COLIEE2023. All methods are assessed under a one stage setting using seven retrieval metrics.
Table 10: Effectiveness of graph contrastive learning. (%)
Methods COLIEE2022
P@5 R@5 Mi-F1 Ma-F1 MRR@5 MAP NDCG@5
w/o GCL 35.9 $\pm$0.5 42.7 $\pm$0.6 39.2 $\pm$0.7 43.3 $\pm$0.5 67.2 $\pm$0.9 64.0 $\pm$1.2 69.4 $\pm$1.2
w/ FeatMask-Node 41.0 $\pm$0.6 48.7 $\pm$0.8 44.6 $\pm$0.7 48.9 $\pm$0.9 74.2 $\pm$1.4 71.4 $\pm$1.2 76.5 $\pm$1.4
w/ FeatMask-Edge 41.1 $\pm$1.1 48.8 $\pm$1.3 44.6 $\pm$1.1 49.2 $\pm$1.3 73.9 $\pm$1.4 71.0 $\pm$1.3 76.6 $\pm$1.1
w/ EdgeDrop 43.7 $\pm$0.2 51.9 $\pm$0.2 47.5 $\pm$0.2 52.4 $\pm$0.3 77.0 $\pm$0.5 74.7 $\pm$0.9 79.3 $\pm$0.8
Methods COLIEE2023
P@5 R@5 Mi-F1 Ma-F1 MRR@5 MAP NDCG@5
w/o GCL 18.0 $\pm$0.3 33.4 $\pm$0.5 23.4 $\pm$0.3 24.0 $\pm$0.2 39.3 $\pm$0.7 38.1 $\pm$0.7 43.2 $\pm$0.5
w/ FeatMask-Node 22.2 $\pm$0.3 41.3 $\pm$0.5 28.9 $\pm$0.4 30.0 $\pm$0.5 48.0 $\pm$0.5 46.2 $\pm$1.0 52.0 $\pm$0.9
w/ FeatMask-Edge 22.6 $\pm$0.3 41.9 $\pm$0.6 29.3 $\pm$0.4 30.4 $\pm$0.5 48.2 $\pm$1.1 46.5 $\pm$1.0 52.3 $\pm$1.0
w/ EdgeDrop 22.6 $\pm$0.4 41.9 $\pm$0.8 29.3 $\pm$0.6 30.5 $\pm$0.7 48.7 $\pm$0.7 47.0 $\pm$0.7 52.6 $\pm$0.8
As shown in Table 10, applying GCL with any augmentation strategy yields substantial improvements over the model trained without GCL on both datasets, confirming that contrastive learning enhances the model’s ability to capture structural and semantic patterns in legal case graphs. Among the three augmentation methods, edge dropping consistently achieves the best performance, outperforming both feature masking strategies across all metrics and datasets. Unlike node or edge feature masking, which alters only the attribute space, edge dropping modifies the structural topology of the graph. This leads to larger perturbations between augmented views, enabling the contrastive objective to learn more discriminative and robust representations. Between the two feature masking augmentations, edge feature masking performs better than node feature masking on both datasets. Since edges in the case graph encode relations between legal entities which are central to legal reasoning and retrieval, masking edge attributes encourages the model to better capture and reconstruct relation semantics. This also aligns with the design of LEXA, which places emphasis on updating and utilizing edge features. The relative improvement of edge masking therefore highlights the importance of rich and well modelled edge representations for achieving higher retrieval accuracy.
5.8 Parameter Sensitivity (RQ7)
In this experiment, the temperature coefficient $\tau$ and the number of easy negative samples in the contrastive loss in Equation (18) are investigated for their parameter sensitivity in LEXA. Specifically, the result of the temperature coefficient $\tau$ is presented in Fig. 2, and the results of the number of easy negative samples are presented in Fig. 3.
Figure 2: Parameter sensitivity for the temperature $\tau$ in the contrastive loss of LEXA.
5.8.1 Temperature coefficient.
As illustrated in Fig. 2, we evaluate the temperature parameter $\tau$ over the range of {0.05,0.1,0.5,1,3}. Across both COLIEE2022 and COLIEE2023, $\tau$=0.1 consistently yields the best Macro-F1 and NDCG results, indicating that a moderate temperature produces the most informative similarity distribution for contrastive learning. Larger temperatures (such as $\tau$=1 or $\tau$=3) reduce performance by overly smoothing the similarity scores, thereby diminishing the discriminative strength of the contrastive objective. In contrast, very small temperatures of $\tau$=0.05 make the similarity distribution too sharp, causing the model to over emphasize a few highly similar pairs and neglect useful relational structure among other samples.
The results above demonstrate that the temperature parameter critically regulates the balance between smoothness and sharpness in the similarity space, with $\tau=0.1$ offering the most stable and effective configuration for contrastive learning in LEXA.
Figure 3: Parameter sensitivity for the number of easy negative samples in the contrastive loss of LEXA.
5.8.2 Number of easy negative samples
As shown in Fig. 3, we investigate the effect of varying the number of easy negative samples using the set {0, 1, 5, 10}. Easy negatives are randomly selected non-relevant cases that help shape the contrastive learning signal. Across both datasets, introducing a small number of one easy negative sample leads to the best overall performance on Ma-F1@5 and NDCG@5, indicating that a limited amount of easy negative information strengthens the contrastive objective without overwhelming the model. When no easy negatives are included, the performance on both datasets is clearly weaker, suggesting that the model benefits from the additional contrast provided by even a minimal set of easy negatives. However, using too many easy negatives (such as 5 or 10) results in a gradual decline in performance. This occurs because the learning objective becomes dominated by trivial negative instances, reducing the relative importance of informative positive and negative distinctions and weakening the supervision signal.
Overall, the results demonstrate that easy negatives are beneficial when used in moderation, and that a small number of such samples provides the most effective balance for training LEXA under the contrastive learning framework.
6 Conclusion
This paper introduced LEXA, an enhanced legal case retrieval framework that extends CaseGNN by fully utilising structural information in text-attributed case graphs. LEXA integrates three key improvements: an edge-updated graph attention layer (EUGAT) for jointly refining node and edge features, an enhanced graph contrastive learning objective for stronger supervision, and LLM-based contextualised embeddings to enrich graph representations. Experiments on COLIEE 2022 and COLIEE 2023 show that LEXA consistently outperforms CaseGNN and achieves state-of-the-art results. These findings highlight the importance of combining structural legal knowledge with contextual LLM representations to advance legal case retrieval.

### v3

@@
 Besides of the graph structure augmentation such as removing nodes and edges, graph feature masking [BGRL, GraphCL] is also a popular graph augmentation method to perform efficient graph augmenting or corrupting by randomly mask the features of nodes or edges.
-Considering graph can be divided into subgraph, another graph augmentation approach called subgraph cropping is to crop out subgraph from the original graph then use the subgraph itself or the remaining graph as samples for contrastive learning (e.g.,  [GraphCL, GraphCrop]).
+Considering graph can be divided into subgraph, another graph augmentation approach called subgraph cropping is to crop out subgraph from the original graph then use the subgraph itself or the remaining graph as samples for contrastive learning (e.g.,  [GraphCL, GraphCrop, puma, tnt-ood]).
 2.3 LLM Embeddings
@@
 Additionally, Qwen3-Embedding-8B[qwen3embedding], trained using an innovative multi-stage pipeline that combines large-scale unsupervised pretraining with supervised fine-tuning on high quality datasets, demonstrates strong performance across both general purpose and domain-specific retrieval tasks.
+Specifically, some LLM-embedding–based legal retrieval methods, such as UQLegalAI@COLIEE2025 [uqlegalai], and legally fine-tuned large language models, such as ReaKase-8B [reakase], have been proposed to enhance semantic understanding of legal texts.
 These advancements highlight the growing effectiveness and specialization of LLM-based embedding models in both general and domain-specific retrieval scenarios, particularly within the legal domain.
@@
 Finally, a contrastive learning objective, detailed in Section 4.5, aligns embeddings of augmented views while preserving discriminative information, resulting in robust and semantically rich legal case representations.
-Figure 1: The framework of LEXA.
-Given legal case $c$, the legal fact and the legal issue sections are converted into case graph.
-The case graph is processed by L layers of EUGAT and a Readout function to obtain an overall case graph representation.
-The whole framework is trained with the contrastive loss with positive, negative and augmented samples.
+Figure 1: Given a legal case $c$, the legal fact and legal issue sections are extracted and transformed into structured case graphs using LEXA-8B and case graph generation (TACG).
+Each case graph is encoded by L stacked Edge-Updated Graph Attention Network (EUGAT) layers to model semantic and relational dependencies.
+A readout function aggregates node representations into graph-level embeddings, which are combined to form the final case representation $h_{c}$.
+The model is trained end-to-end with a contrastive learning objective incorporating positive, easy and hard negative, and augmented samples to improve discriminative power and robustness.
 4.1 LEXA-8B
@@
 In Equation 18, the augmented graph of relevant cases $d^{+}$ is denoted as $d_{aug}^{+}$ while the augmented graph of easy negative sample $d^{easy-}$ is denoted as $d_{aug}^{easy-}$.
-During the graph contrastive learning process, only positive and easy negative samples are augmented; hard negative samples $d^{hard-}$ are used in their original form without augmentation.
-This design is motivated by experimental observations indicating that a limited number of hard negative samples is sufficient for effective training, whereas an excessive number can lead to notable performance degradation.
-Furthermore, for each case graph, only a single augmentation strategy is employed to generate an augmented graph.
+For each case graph, only a single augmentation strategy is employed to generate an augmented graph.
 This decision is motivated by the empirical finding that using multiple augmentation methods simultaneously results in reduced performance.
 Additionally, during training, either augmented positive samples or augmented easy negative samples are used individually, as combining both was found to yield suboptimal results.
-The complete model pipeline is illustrated in Fig.
-1.
 5 Experiments
@@
   • RQ3: How effective is EUGAT compared with other GNNs?
-  • RQ4: How effective is prompt-based LLM features encoding?
-  • RQ5: How does graph contrastive learning help with LCR in LEXA?
-  • RQ6: How do hyper-parameter settings affect LEXA?
+  • RQ4: How different LLMs affect the performance
+  • RQ5: How effective is prompt-based LLM features encoding?
+  • RQ6: How does graph contrastive learning help with LCR in LEXA?
+  • RQ7: How does graph sparsity affect retrieval performance in LEXA?
+  • RQ8: How do hyper-parameter settings affect LEXA?
 5.1 Setup
@@
 BM25 17.9 21.2 19.4 21.4 23.6 25.4 33.6
-LM
-LEGAL-BERT 4.47 5.30 4.85 5.38 7.42 7.47 10.9
+LM LEGAL-BERT 4.47 5.30 4.85 5.38 7.42 7.47 10.9
 MonoT5 0.71 0.65 0.60 0.79 1.39 1.41 1.73
@@
 PromptCase 17.1 20.3 18.5 20.5 35.1 33.9 38.7
-LLM
-E5-Mistral-7B-Instruct 21.4 25.4 23.2 25.7 26.8 28.5 38.0
+LLM E5-Mistral-7B-Instruct 21.4 25.4 23.2 25.7 26.8 28.5 38.0
 Qwen3-Embedding-8B 21.6 25.7 23.5 26.0 26.7 29.0 37.8
 Inf-Retriever-V1 21.1 25.1 22.9 25.5 26.6 28.7 37.9
-Graph
-CaseGNN 35.5 $\pm$0.2 42.1 $\pm$0.2 38.4 $\pm$0.3 42.4 $\pm$0.1 66.8 $\pm$0.8 64.4 $\pm$0.9 69.3 $\pm$0.8
+Graph CaseGNN 35.5 $\pm$0.2 42.1 $\pm$0.2 38.4 $\pm$0.3 42.4 $\pm$0.1 66.8 $\pm$0.8 64.4 $\pm$0.9 69.3 $\pm$0.8
 CaseLink 37.0 $\pm$0.1 43.9 $\pm$0.1 40.1 $\pm$0.1 44.2 $\pm$0.1 67.3 $\pm$0.5 65.0 $\pm$0.2 70.3 $\pm$0.1
@@
 Two-stage
-LM
-SAILER 23.8 25.7 24.7 25.2 43.9 42.7 48.4
+LM SAILER 23.8 25.7 24.7 25.2 43.9 42.7 48.4
 PromptCase 23.5 25.3 24.4 30.3 41.2 39.6 45.1
-LLM
-E5-Mistral-7B-Instruct 22.0 26.2 24.0 25.8 48.4 46.8 51.5
+LLM E5-Mistral-7B-Instruct 22.0 26.2 24.0 25.8 48.4 46.8 51.5
 Qwen3-Embedding-8B 21.9 26.0 23.8 25.8 48.8 48.0 52.1
 Inf-Retriever-V1 22.2 26.4 24.1 26.1 45.5 44.5 49.6
-Graph
-CaseGNN 22.9 $\pm$0.1 27.2 $\pm$0.1 24.9 $\pm$0.1 27.0 $\pm$0.1 54.9 $\pm$0.4 54.0 $\pm$0.5 57.3 $\pm$0.6
+Graph CaseGNN 22.9 $\pm$0.1 27.2 $\pm$0.1 24.9 $\pm$0.1 27.0 $\pm$0.1 54.9 $\pm$0.4 54.0 $\pm$0.5 57.3 $\pm$0.6
 CaseLink 24.7 $\pm$0.1 29.1 $\pm$0.1 26.8 $\pm$0.1 29.2 $\pm$0.1 56.0 $\pm$0.2 55.0 $\pm$0.2 58.6 $\pm$0.1
@@
 BM25 16.5 30.6 21.4 22.2 23.1 20.4 23.7
-LM
-LEGAL-BERT 4.64 8.61 6.03 6.03 11.4 11.3 13.6
+LM LEGAL-BERT 4.64 8.61 6.03 6.03 11.4 11.3 13.6
 MonoT5 0.38 0.70 0.49 0.47 1.17 1.33 0.61
@@
 PromptCase 16.0 29.7 20.8 21.5 32.7 32.0 36.2
-LLM
-E5-Mistral-7b-Instruct 16.0 29.7 20.8 21.5 21.9 22.9 30.8
+LLM E5-Mistral-7b-Instruct 16.0 29.7 20.8 21.5 21.9 22.9 30.8
 Qwen3-Embedding-8B 18.4 34.1 23.9 25.2 25.8 27.5 36.4
 Inf-Retriever-V1 19.0 35.3 24.7 26.0 26.5 28.0 37.6
-Graph
-CaseGNN 17.7 $\pm$0.7 32.8 $\pm$0.7 23.0 $\pm$0.5 23.6 $\pm$0.5 38.9 $\pm$1.1 37.7 $\pm$0.8 42.8 $\pm$0.7
+Graph CaseGNN 17.7 $\pm$0.7 32.8 $\pm$0.7 23.0 $\pm$0.5 23.6 $\pm$0.5 38.9 $\pm$1.1 37.7 $\pm$0.8 42.8 $\pm$0.7
 CaseLink 20.9 $\pm$0.3 38.4 $\pm$0.6 27.1 $\pm$0.3 28.2 $\pm$0.3 45.8 $\pm$0.5 44.3 $\pm$0.7 49.8 $\pm$0.4
@@
 Two-stage
-LM
-SAILER 19.6 32.6 24.5 23.5 37.3 36.1 40.8
+LM SAILER 19.6 32.6 24.5 23.5 37.3 36.1 40.8
 PromptCase 21.8 36.3 27.2 26.5 39.9 38.7 44.0
-LLM
-E5-Mistral-7b-Instruct 19.7 36.6 25.6 26.5 46.9 44.4 49.9
+LLM E5-Mistral-7b-Instruct 19.7 36.6 25.6 26.5 46.9 44.4 49.9
 Qwen3-Embedding-8B 20.1 37.4 26.2 27.0 47.7 46.3 51.1
 Inf-Retriever-V1 19.7 36.6 25.6 26.6 48.2 46.9 51.6
-Graph
-CaseGNN 20.2 $\pm$0.2 37.6 $\pm$0.5 26.3 $\pm$0.3 27.3 $\pm$0.2 45.8 $\pm$0.9 44.4 $\pm$0.8 49.6 $\pm$0.8
+Graph CaseGNN 20.2 $\pm$0.2 37.6 $\pm$0.5 26.3 $\pm$0.3 27.3 $\pm$0.2 45.8 $\pm$0.9 44.4 $\pm$0.8 49.6 $\pm$0.8
 CaseLink 21.0 $\pm$0.3 38.9 $\pm$0.5 27.1 $\pm$0.3 28.2 $\pm$0.3 48.8 $\pm$0.2 47.2 $\pm$0.1 52.6 $\pm$0.1
@@
 LEXA 22.6 $\pm$0.4 41.9 $\pm$0.8 29.3 $\pm$0.6 30.5 $\pm$0.7 48.7 $\pm$0.7 47.0 $\pm$0.7 52.6 $\pm$0.8
-In this section, the effectiveness of various LLMs for feature encoding is evaluated.
+In this section, the effectiveness of various LLMs for feature encoding is evaluated on both datasets across all evaluation metrics, with results reported in Table 7.
 Both E5-Mistral-7B-Instruct and Qwen3-Embedding-8B are employed to encode relation triplets into node and edge feature embeddings.
@@
 The relative improvement of edge masking therefore highlights the importance of rich and well modelled edge representations for achieving higher retrieval accuracy.
-5.8 Parameter Sensitivity (RQ7)
+5.8 Impact of Graph Sparsification via Edge Pruning (RQ7)
+Table 11: Impact of graph sparsification via edge pruning.
+Methods COLIEE2022
+P@5 R@5 Mi-F1 Ma-F1 MRR@5 MAP NDCG@5
+80% Edge 39.6 $\pm$0.4 47.1 $\pm$0.5 43.0 $\pm$0.5 47.7 $\pm$0.5 71.0 $\pm$0.3 68.4 $\pm$0.3 74.5 $\pm$0.1
+90% Edge 40.5 $\pm$0.5 48.1 $\pm$0.6 43.9 $\pm$0.6 48.6 $\pm$0.7 72.6 $\pm$2.2 70.0 $\pm$1.6 75.7 $\pm$1.5
+100% Edge 43.7 $\pm$0.2 51.9 $\pm$0.2 47.5 $\pm$0.2 52.4 $\pm$0.3 77.0 $\pm$0.5 74.7 $\pm$0.9 79.3 $\pm$0.8
+Methods COLIEE2023
+P@5 R@5 Mi-F1 Ma-F1 MRR@5 MAP NDCG@5
+80% Edge 22.2 $\pm$0.1 41.3 $\pm$0.1 28.9 $\pm$0.1 29.9 $\pm$0.1 47.9 $\pm$0.3 46.1 $\pm$0.3 51.7 $\pm$0.3
+90% Edge 22.4 $\pm$0.4 41.7 $\pm$0.2 29.2 $\pm$0.2 30.3 $\pm$0.2 48.6 $\pm$0.4 46.8 $\pm$0.6 51.4 $\pm$0.5
+100% Edge 22.6 $\pm$0.4 41.9 $\pm$0.8 29.3 $\pm$0.6 30.5 $\pm$0.7 48.7 $\pm$0.7 47.0 $\pm$0.7 52.6 $\pm$0.8
+To evaluate how reducing graph density affects retrieval effectiveness, three experiments with different edge pruning ratios are conducted on both datasets, as shown in Table 11.
+Specifically, the 80% edge setting retains 80% of the edges from the original TACG graph, the 90% edge setting retains 90% of the edges, and the 100% edge setting corresponds to the original, unpruned graph.
+Overall, the results indicate that reducing graph density through edge pruning leads to a steady degradation in retrieval effectiveness across all evaluation metrics.
+The unpruned graph consistently achieves the best performance on both datasets, outperforming pruned variants in terms of P@5, F1, MAP, MRR@5, and NDCG@5.
+Moderate edge pruning (90%) results in noticeable but limited performance drops, while more aggressive pruning (80%) causes further degradation.
+This trend suggests that excessive edge removal disrupts important structural and semantic relationships that are essential for accurate legal case retrieval.
+Although the performance gap between pruned and unpruned graphs is smaller on COLIEE2023 than on COLIEE2022, the fully connected graph remains consistently superior.
+These findings demonstrate that preserving rich graph connectivity is crucial for stable ranking quality across different dataset scales, and naive edge pruning introduces a clear accuracy–efficiency trade-off.
+5.9 Parameter Sensitivity (RQ8)
 In this experiment, the temperature coefficient $\tau$ and the number of easy negative samples in the contrastive loss in Equation (18) are investigated for their parameter sensitivity in LEXA.
@@
 These findings highlight the importance of combining structural legal knowledge with contextual LLM representations to advance legal case retrieval.
+Declarations
+Funding declaration This work is supported by Australian Research Council CE200100025, DP230101196, DP230101753 and DE250100919.
+Ethics declaration Not applicable.